# Search for Solar Boosted Dark Matter Particles at the PandaX-4T Experiment Guofang Shen

### v1

@@ Abstract
We present a novel constraint on light dark matter utilizing tonneyear of data acquired from the PandaX-4T dual-phase xenon time projection chamber. This constraint is derived through detecting electronic recoil signals resulting from the interaction with solar-enhanced dark matter flux. Low-mass dark matter particles, lighter than a few MeV/, can scatter with the thermal electrons in the Sun. Consequently, with higher kinetic energy, the boosted dark matter component becomes detectable via contact scattering with xenon electrons, resulting in a few keV energy deposition that exceeds the threshold of PandaX-4T. We calculate the expected recoil energy in PandaX-4T considering the Sun’s acceleration and the detection capabilities of the xenon detector. The first experimental search results using the xenon detector yield the most stringent cross-section of at / for a solar boosted dark matter mass ranging from to /, achieving a 23 fold improvement compared with earlier experimental studies.
@@
The freeze-out mechanism proposes that dark matter was formed when particles frequently collided within a dense, high-temperature ”soup” in the early universe. In this model, most particles initially exist as dark matter but later are destroyed. The remaining DM particles are usually classified as weakly interacting massive particles (WIMPs) Steigman et al. (2012). In contrast, freeze-in represents a complementary process in which dark matter arises by gradually replacing ordinary matter Bernal et al. (2017). In recent decades, WIMPs have emerged as the predominant candidates for elucidating the phenomenon of missing mass, notwithstanding the lack of compelling evidence from direct detection experiments utilizing xenon or other materials Undagoitia and Rauch (2015); Liu et al. (2017); Roszkowski et al. (2018); Schumann (2019). In recent years, xenon-based detectors such as PandaX-4T, LZ, and XENONnT have provided the most stringent limits on WIMPs–nucleon spin-independent interactions, pushing the cross-section down to for a DM mass of , for , and for , respectively Bo et al. (2024); Aalbers et al. (2024); Aprile et al. (2023).
@@
The traditional method for directly detecting DM is to measure the elastic scattering between target nuclei and DM particles. Since cold DM in the halo is non-relativistic, the direct detection strategy is only sensitive to DM with masses above sub-GeV due to the detector thresholds. In recent years, pushing the search to lower mass scales of sub-GeV or even sub-MeV DM, which are still untapped for promising thermal relic DM possibilities, has spurred many studies. Various approaches have been proposed, such as accessing the electron recoil (ER) channel Essig et al. (2012a); Emken et al. (2019); Essig et al. (2012b, 2016); Agnes et al. (2018a); Amaral et al. (2020); Barak et al. (2020); Aguilar-Arevalo et al. (2019); Cheng et al. (2021a), lowering the nuclear recoil (NR) detection threshold Pirro and Mauskopf (2017); Abdelhameed et al. (2019); Mancuso et al. (2020); Aprile et al. (2019a); Agnes et al. (2018b), and utilizing the so-called Migdal effect Ibe et al. (2018); Baxter et al. (2020); Essig et al. (2020); Akerib et al. (2019); Armengaud et al. (2019); Liu et al. (2019); Aprile et al. (2019b).
@@
We focus exclusively on the generation and detection of SBDM through electron scattering An et al. (2018, 2021), as the contribution of ions to MeV scale SBDM is negligible Emken (2022).
The ambient DM with a typical velocity between and would be attracted by the Sun via gravity when they pass nearby. The tremendous amount of energetic electrons in the Sun with a density at the level of could effectively heat the DM particles and boost their kinetic energy to keV levels An et al. (2021). The SBDM could then escape the Sun and reach terrestrial detectors like PandaX-4T Meng et al. (2021) leaving detectable signals that exceed the threshold with the same amplitude on the bounded electron of the xenon. Calculating the flux of SBDM emanating from the surface of the Sun is imperative. Subsequently, we will assess the event rate in the xenon detector based on the flux and for different DM masses.
@@
where is the DM flux in the local Milky Way halo, and is the scattering radius determined through the radius-temperature relation Bahcall et al. (2005). indicates the gravitational focusing effect, which enhances the solar scatterings and has a value of O() when An et al. (2018). is the Sun-Earth distance. The is the electron number density within the radius of the Sun’s core (). The mean free path for a DM particle inside the core of the Sun is which would dramatically affect the calculations of the flux when .
The possibility of multiple scattering should be considered in Eq. 1 but is difficult to treat analytically. So, a Monte Carlo simulation program An et al. (2021) is built to mimic the propagation of DM particles inside the Sun with an elaborate solar model taken fromSerenelli et al. (2009). The impact parameter is randomly chosen from to to incorporate gravitational focusing. Beyond this range, only a small proportion of the slowest DM could be trapped by the gravity field, minimally contributing to the boosted flux. is the simulated kinetic energy distribution of an SBDM particle, where , the DM mass, and are two key input parameters for the simulation. The boosted flux observed at the detector’s location is illustrated as
where is the area where the DM particles could be boosted in the simulation.
@@
The entire calculation is based on the framework Essig et al. (2017) with the following modifications. First, we use xenon as the target; second, we sum the electron recoiling and deexcitation photon energy from different electron shells in xenon; last,
@@
Next, we calculate the differential recoil rate per unit detector
@@
Fig. 2 illustrates the expected event rate for different DM masses with representative cross- sections and the total efficiencies that combine reconstruction, quality cuts, and region-of-interest (ROI) efficiencies. The total efficiencies are estimated using both the data-driven and waveform-simulation method Li et al. (2024); Luo et al. (2024). Furthermore, the energy resolution curves of Run0 and Run1 are given in Fig. 2. The structure of the energy spectrum is attributed to interactions between the SBDM and different shell electrons of the xenon. The recoil energy could reach four keV, beyond the PandaX-4T experiment threshold Meng et al. (2021); Bo et al. (2024).
Figure 2: The expected event rate of SBDM in the xenon detector with = 0.25 MeV, 0.5 MeV, and 1.0 MeV for . For = 0.5 MeV, the event rate with from to cross-sections is also shown. The dashed green and mint orange lines represent the total efficiencies of Run0 and Run1, and the shield bands illustrate the efficiency uncertainties. The dotted and dashed magenta curves are the energy resolution of Run0 and Run1, respectively.
@@
PDE, EEE, and represent the photon detection efficiency, electron extraction efficiency, and single electron gain, respectively. The value is obtained by measuring the smallest signals Meng et al. (2021). The values of PDE and EEE are predetermined by calibrating the monoenergetic electronic recoil peaks: 41.5 keV (83m Kru), 164 keV (131mXe), 236 keV (129mXe), and 408 keV (127Xe).
his analysis uses the same data set fsrom Run0 and Run1 as Ref. Bo et al. (2024), including 11 subsets segmented according to slightly different electric field configurations and background levels. The fiducial masses are tonne for Run0 and tonne for Run1. The total exposure is tonneyear.
@@
The energy of SBDM signals ranges from to keV. To enhance the measurement of the ER background from 222Rn, we extended the fitting range from 4 to keV as a sideband. Furthermore, a threshold of 99.5% ER acceptance cut was implemented to exclude most NR events arising from solar 8B neutrinos and neutron interactions with the xenon nucleus as well as radon plateout on the surface of the PTFE reflector of the PandaX-4T TPC. This cut is derived from the simulated data of the PandaX-4T signal model Luo et al. (2024). Finally, we consider the primary ER backgrounds, including tritium, 85Kr, 214Pb, 212Pb, 127Xe, 136Xe, 124Xe, neutrino, material, and accidental events. The detailed details on each component are discussed in Zeng et al. (2024).
@@
FDataifrom andaX-4T,1 (Run0) and (Run1)ER events are selected TAprofile likelihood ratio( PLR) approach Baxter (2021) is used in this analysis with a one-dimensional signal energy spectrum based on HistFitter Cranmer et al. (2012). In this analysis, both background-only and background-plus-signal hypotheses are tested. Background contributions are well aligned with the expected values, and all nuisance parameters remain within of the input values. The best fit of the SBDM result is illustrated in Fig. 3 and the expected events from each component are summarized in Table 1. Ultimately, no significant SBDM signal is detected above the background.
Figure 3: Best fit results of Run0 and Run1 data with background-plus-signal hypothesis. Error bars are statistical uncertainties.
@@
For comparison, recent results from the phenomenological study An et al. (2018), the SBDM analysis utilizing the HPGe detector from CDEX CDEXSRDM, and the constraints derived from cosmological and astrophysical observables related to freeze-out DM Planck Collaboration et al. (2016), and other analyses of sub-GeV DM PandaX-4TS2only; Cheng et al. (2021b); Aprile et al. (2019c); Essig et al. (2012b) are overlaid. Our results cover a large region from 20 keV/
@@
. The lowest 90% confidence level on the SBDM cross-section is for an SBDM mass of . Our results provide the most stringent constraints, achieving a 23-fold improvement compared to recent experimental studies conducted by CDEX CDEXSRDM.
Figure 4: 90% confidence level (C.L.) exclusion upper limits (red line) and exclusion region (lightgreen region) for the DM-electron cross-section are presented for SBDM. The light gray band is stellar cooling constraints from red giant (RG) stars Chang et al. (2021). The gray curve denotes the freeze-out DM mechanism Planck Collaboration et al. (2016), while the dashed yellow line indicates the SBDM limits from Ref. An et al. (2018). Additionally, the orange line is from the CDEX experimental result CDEXSRDM, and the direct constraints based on S2 analysis from Ref. PandaX-4TS2only; Cheng et al. (2021b); Aprile et al. (2019c); Essig et al. (2012b) are also overlaid at the right side.
@@
Sun with the 1.54 tonneyear data of Run0 and Run1 from PandaX-4T. This leads to a new sensitivity at the sub-pb level for light DM in the sub-MeV mass range. No significant dark matter signals are identified above the expected background. A new upper limit exclusion is established for the SBDM-electron interactions, robustly excluding sub-MeV DM with a scattering cross-section with electrons within to . The well-known freeze-out mechanism below the region could be excluded. Combined with analyses such as ionization signals PandaX-4TS2only, future upgraded detectors will test the small open window between .
@@
Szydagis et al. (2011)M. Szydagis, N. Barry, K. Kazkaz, J. Mock, D. Stolp, M. Sweany, M. Tripathi, S. Uvarov, N. Walsh, and M. Woods, Journal of Instrumentation 6, P10002 (2011).

### v2

@@ Abstract
We present a novel constraint on light dark matter utilizing tonneyear of data acquired from the PandaX-4T dual-phase xenon time projection chamber. This constraint is derived through detecting electronic recoil signals resulting from the interaction with solar-enhanced dark matter flux. Low-mass dark matter particles, lighter than a few MeV/, can scatter with the thermal electrons in the Sun. Consequently, with higher kinetic energy, the boosted dark matter component becomes detectable via contact scattering with xenon electrons, resulting in a few keV energy deposition that exceeds the threshold of PandaX-4T. We calculate the expected recoil energy in PandaX-4T considering the Sun’s acceleration with heavy mediators and the detection capabilities of the xenon detector. The first experimental search results using the xenon detector yield the most stringent upper limits cross-section of at / for a solar boosted dark matter mass ranging from to /, achieving a 23 fold improvement compared with earlier experimental studies.
@@
The freeze-out mechanism proposes that dark matter was formed when particles frequently collided within a dense, high-temperature ”soup” in the early universe. In this model, most particles initially exist as dark matter but later are destroyed. The remaining DM particles are usually classified as weakly interacting massive particles (WIMPs) Steigman et al. (2012).
In recent decades, WIMPs have emerged as the predominant candidates for elucidating the phenomenon of missing mass, notwithstanding the lack of compelling evidence from direct detection experiments utilizing xenon or other materials Undagoitia and Rauch (2015); Liu et al. (2017); Roszkowski et al. (2018); Schumann (2019). In recent years, xenon-based detectors such as PandaX-4T, LZ, and XENONnT have provided the most stringent limits on WIMPs–nucleon spin-independent interactions, pushing the cross-section down to for a DM mass of , for , and for , respectively Bo et al. (2024); Aalbers et al. (2024); Aprile et al. (2023).
@@
The traditional method for directly detecting DM is to measure the elastic scattering between target nuclei and DM particles. Since cold DM in the halo is non-relativistic, the direct detection strategy is only sensitive to DM with masses exceeding sub-GeV due to the detector thresholds. In recent years, pushing the search to lower mass scales of sub-GeV or even sub-MeV DM, which are still untapped for promising thermal relic DM possibilities, has spurred many studies. Various approaches have been proposed, such as accessing the electron recoil (ER) channel Essig et al. (2012a); Emken et al. (2019); Essig et al. (2012b, 2016); Agnes et al. (2018a); Amaral et al. (2020); Barak et al. (2020); Aguilar-Arevalo et al. (2019); Cheng et al. (2021a), lowering the nuclear recoil (NR) detection threshold Pirro and Mauskopf (2017); Abdelhameed et al. (2019); Mancuso et al. (2020); Aprile et al. (2019a); Agnes et al. (2018b), and utilizing the so-called Migdal effect Ibe et al. (2018); Baxter et al. (2020); Essig et al. (2020); Akerib et al. (2019); Armengaud et al. (2019); Liu et al. (2019); Aprile et al. (2019b).
@@
In this letter, we focus exclusively on the generation and detection of SBDM through electron scattering An et al. (2018, 2021), as the contribution of ions to MeV scale SBDM is negligible Emken (2022).
The ambient DM with a typical velocity between and would be attracted by the Sun via gravity when they pass nearby. The tremendous amount of energetic electrons in the Sun with a density at the level of could effectively heat the DM particles and boost their kinetic energy to keV levels An et al. (2021). The SBDM could then escape the Sun and reach terrestrial detectors like PandaX-4T Meng et al. (2021) leaving detectable signals that exceed the threshold with the same amplitude on the bounded electrons of the xenon. Calculating the flux of SBDM emanating from the surface of the Sun is imperative. Subsequently, we will assess the event rate in the xenon detector based on the flux and for different DM masses.
@@
where is the DM flux in the local Milky Way halo with a DM local density Catena and Ullio (2010) and a typical DM velocity distribution, and is the scattering radius determined through the radius-temperature relation Bahcall et al. (2005). indicates the gravitational focusing effect, which enhances the solar scatterings and has a value of O() when An et al. (2018). is the Sun-Earth distance. The is the electron number density within the radius of the Sun’s core (). The mean free path for a DM particle inside the core of the Sun is which would dramatically affect the calculations of the flux when .
The possibility of multiple scattering should be considered in Eq. 1 but is difficult to treat analytically. So, a Monte Carlo simulation program An et al. (2018) is built to mimic the propagation of DM particles inside the Sun with an elaborate solar model taken from Serenelli et al. (2009).
In the simulation, the perpendicular distance of halo DM particles to the Sun, which is known as impact parameter , is confined within to incorporate gravitational focusing in the calculation of the SBDM. Beyond this range, only a small proportion of the slowest DM could be trapped by the gravity field, minimally contributing to the boosted flux. is the simulated kinetic energy distribution of an SBDM particle, where , the DM mass, and are two key input parameters for the simulation. The boosted flux observed at the detector’s location is illustrated as
where is the area where the DM particles could be boosted in the simulation and .
@@
There is a framework Essig et al. (2016) built to calculate sub-GeV dark matter scattering off electrons with
semiconductor targets.
Our entire calculation is based on the framework to calculate the DM-electron scattering rate, but with the following modifications.
First, we use xenon as the target; second, we sum the electron recoiling and deexcitation photon energy from different electron shells in xenon; last,
@@
Next, we calculate the differential recoil rate
@@
PDE, EEE, and represent the photon detection efficiency, electron extraction efficiency, and single electron gain, respectively. The here already involve the PDE for S2 and are obtained by measuring the smallest recorded signals Meng et al. (2021). The values of PDE and EEE are predetermined by calibrating the monoenergetic electronic recoil peaks: 41.5 keV (83m Kr), 164 keV (131mXe), 236 keV (129mXe), and 408 keV (127Xe).
Fig. 2 illustrates the expected interaction rate for different DM masses with representative cross-sections and the total efficiencies in PandaX-4T. The total efficiencies are estimated using both the data-driven and waveform-simulation method Li et al. (2024); Luo et al. (2024). Furthermore, the energy resolution curves of Run0 and Run1 are given in Fig. 2. The structure of the energy spectrum is attributed to interactions between the SBDM and different shell electrons of the xenon. The recoil energy could reach four keV, beyond the PandaX-4T experiment threshold Meng et al. (2021); Bo et al. (2024).
Figure 2: The expected interaction rate of SBDM in the xenon detector with = 0.25 MeV, 0.5 MeV, and 1.0 MeV for . For = 0.5 MeV, the expected interaction rate with from to cross-sections is also shown. The dashed green and dashed orange lines represent the total efficiencies of Run0 and Run1, and the shield bands illustrate the efficiency uncertainties. The dotted and dashed magenta curves are the energy resolution of Run0 and Run1, respectively. With interaction rate, efficiency and energy resolution convolved, the corresponding reconstructed signal spectrum could be derived and used in the following analysis.
This analysis uses the same data sets from Run0 and Run1 as Ref. Bo et al. (2024), including 11 subsets segmented according to slightly different electric field configurations and background levels. The fiducial masses are tonne for Run0 and tonne for Run1. The total exposure is tonneyear.
@@
The energy of SBDM signals ranges from to keV. To enhance the measurement of the ER background from 222Rn, we choose the final region-of-interest (ROI) from 0 to 30 keV. Furthermore, a threshold of 99.5% ER acceptance cut was implemented to exclude most NR events arising from solar 8B neutrinos and neutron interactions with the xenon nucleus as well as radon plateout on the surface of the PTFE reflector of the PandaX-4T TPC. This cut is derived from the simulated data of the PandaX-4T signal model Luo et al. (2024). Finally, we consider the primary ER backgrounds, including tritium, 85Kr, 214Pb, 212Pb, 127Xe, 136Xe, 124Xe, neutrino, material, and accidental events. A comprehensive examination of each component is presented in Zeng et al. (2024).
@@
Data from the PandaX-4T experiments, specifically the
1197 (Run0) and 1431 (Run1) electron recoil (ER) events, have been meticulously selected for this analysis. A profile likelihood ratio (PLR) approach Baxter (2021) is used in this analysis with a one-dimensional signal energy spectrum based on HistFitter Baak et al. (2015). In this analysis, both background-only and background-plus-signal hypotheses are tested. Background contributions are well aligned with the expected values, and all nuisance parameters remain within of the input values. The best fit of the SBDM result is illustrated in Fig. 3 and the expected events from each component are summarized in Table 1. Ultimately, no significant SBDM signal is detected above the background.
Figure 3: Best combined fit results of Run0 and Run1 data with background-plus-signal hypothesis. Error bars are statistical uncertainties.
@@
For comparison, recent results from the phenomenological study An et al. (2018), the SBDM analysis utilizing the HPGe detector from CDEX Zhang et al. (2024), and the constraints derived from cosmological and astrophysical observables related to freeze-out DM Planck Collaboration et al. (2016), and other analyses of sub-GeV DM Li et al. (2023); Cheng et al. (2021b); Aprile et al. (2019c); Essig et al. (2012b) are overlaid. Our results cover a large region from 20 keV/
@@
. The lowest 90% confidence level on the SBDM cross-section is for an SBDM mass of . Our results provide the most stringent constraints, achieving a 23-fold improvement compared to recent experimental studies conducted by CDEX Zhang et al. (2024).
Figure 4: 90% confidence level (C.L.) exclusion upper limits (red line) and exclusion region (lightgreen region) for the DM-electron cross-section are presented for SBDM. The light gray band is stellar cooling constraints from red giant (RG) stars Chang et al. (2021). The gray curve denotes the freeze-out DM mechanism Planck Collaboration et al. (2016), while the dashed yellow line indicates the SBDM limits from Ref. An et al. (2018). Additionally, the orange line is from the CDEX experimental result Zhang et al. (2024), and the direct constraints based on S2 analysis from Ref. Li et al. (2023); Cheng et al. (2021b); Aprile et al. (2019c); Essig et al. (2012b) are also overlaid at the right side.
@@
Sun with the 1.54 tonneyear data of Run0 and Run1 from PandaX-4T. This leads to a new sensitivity at the sub-pb level for light DM in the sub-MeV mass range. No significant dark matter signals are identified above the expected background. A new upper limit exclusion is established for the SBDM-electron interactions with heavy mediators, robustly excluding sub-MeV DM with a scattering cross-section with electrons within to . The well-known freeze-out mechanism below the region could be excluded. Combined with analyses such as ionization signals Li et al. (2023),
The future upgraded PandaX-4T detectors, with significantly enhanced sensitivity and optimized S2-only event detection thresholds Abdukerim et al. (2024), will explore the narrow unconstrained parameter window in the mass range of .
@@
Szydagis et al. (2011)M. Szydagis, N. Barry,
K. Kazkaz, J. Mock, D. Stolp, M. Sweany, M. Tripathi, S. Uvarov, N. Walsh, and M. Woods, Journal of Instrumentation 6, P10002 (2011).